# Unified Multimodal Understanding and Generation Models: Advances, Challenges, and Opportunities

### v5

@@ 3.3 Fused Autoregressive and Diffusion Models
1) Pixel-based Encoding: As shown in Fig. 5 (c-1), pixel-based encoding transforms images into either discrete tokens or continuous latent vectors, which are then used as targets in a diffusion-based denoising process conditioned on autoregressively generated text tokens. Among recent works, Transfusion [38], MonoFormer [37], and LMFusion [188] all adopt continuous latent representations extracted via SD-VAE. These models share a common training objective that combines autoregressive loss for language modeling and diffusion loss for image reconstruction, and utilize bidirectional attention to enable spatial coherence. Despite this shared framework, each model introduces distinct architectural innovations: Transfusion [38] proposes a unified transformer backbone with modality-specific layers to jointly handle discrete and continuous inputs; MonoFormer [37] introduces a compact architecture with shared blocks and task-dependent attention masking to balance AR and diffusion tasks; and LMFusion [188] enables frozen LLMs to perform high-quality image generation through a lightweight visual injection module, preserving language capabilities while training only the vision branch. In contrast, Show-o [39] employs a discrete pixel-based tokenizer based on MAGVIT-v2 [213], generating symbolic image tokens compatible with transformer-style decoding. It supports both AR-based text token generation and diffusion-based image synthesis, supervised through a combination of autoregressive and diffusion losses. Collectively, these models demonstrate the effectiveness of pixel-based encoding in balancing semantic controllability from language models and high-resolution visual fidelity from diffusion processes.
@@
2) Hybrid Encoding: As shown in Fig. 5 (c-2), hybrid encoding fuses both semantic features (e.g., from CLIP or ViT encoders) and pixel-level latents (e.g., from SD-VAE), providing a more expressive image representation. This approach allows models to leverage high-level semantic abstraction while maintaining detailed visual information. Specifically, Janus-flow [189], Mogao [190] and BAGEL [191] adopt a dual-encoder architecture and presents a minimalist architecture that harmonizes AR language models with rectified flow. They decouples the understanding and generation encoders, using SigLIP or the concatenation of SigLIP and SDXL-VAE as the vision encoder for multimodal understanding and SDXL-VAE or FLUX-VAE for image generation. However, the pseudo hybrid encoding design limits the model’s ability to simultaneously leverage both semantic and pixel-level features during generation, as only the pixel encoder is active in the image synthesis process. This decoupling, while beneficial for modularity and training efficiency, prevents the model from fully exploiting semantic cues during image decoding, potentially weakening fine-grained alignment and multimodal compositionality in generative tasks.
@@ 3.4 Any-to-Any Multimodal Models
A recent and notable addition to the any-to-any unified multimodal models is M2-omni [200], which introduces a highly versatile architecture capable of processing and generating a wide variety of modalities, including text, image, video, and audio. M2-omini takes a step forward by incorporating multiple modality-specific tokenizers and decoders, each carefully designed to handle the unique characteristics of different data types. Specifically, it utilizes NaViT [235] to encode videos and images of arbitrary resolution, and combines a pre-trained SD-3 [226] as the image decoder. For audio, M2-omini introduces paraformer-zh [236] to extract audio tokens, and feeds the predicted discrete audio tokens into the pretrained CosyVoice [237] flow matching and vocoder model to generate audio streams. This integration ensures that M2-omini can effectively generate high-quality images, and audio streams from various inputs, making it a truly multi-modal powerhouse.

### v6

@@ 3.3 Fused Autoregressive and Diffusion Models
1) Pixel-based Encoding: As shown in Fig. 5 (c-1), pixel-based encoding transforms images into either discrete tokens or continuous latent vectors, which are then used as targets in a diffusion-based denoising process conditioned on autoregressively generated text tokens. Among recent works, Transfusion [391], MonoFormer [383], and LMFusion [256] all adopt continuous latent representations extracted via SD-VAE. These models share a common training objective that combines autoregressive loss for language modeling and diffusion loss for image reconstruction, and utilize bidirectional attention to enable spatial coherence. Despite this shared framework, each model introduces distinct architectural innovations: Transfusion [391] proposes a unified transformer backbone with modality-specific layers to jointly handle discrete and continuous inputs; MonoFormer [383] introduces a compact architecture with shared blocks and task-dependent attention masking to balance AR and diffusion tasks; and LMFusion [256] enables frozen LLMs to perform high-quality image generation through a lightweight visual injection module, preserving language capabilities while training only the vision branch. In contrast, Show-o [334] employs a discrete pixel-based tokenizer based on MAGVIT-v2 [358], generating symbolic image tokens compatible with transformer-style decoding. It supports both AR-based text token generation and diffusion-based image synthesis, supervised through a combination of autoregressive and diffusion losses. TUNA [194] achieves unified visual representations via a VAE encoder followed with a semantic encoder. Collectively, these models demonstrate the effectiveness of pixel-based encoding in balancing semantic controllability from language models and high-resolution visual fidelity from diffusion processes.
@@
2) Hybrid Encoding: As shown in Fig. 5 (c-2), hybrid encoding fuses both semantic features (e.g., from CLIP or ViT encoders) and pixel-level latents (e.g., from SD-VAE), providing a more expressive image representation. This approach allows models to leverage high-level semantic abstraction while maintaining detailed visual information. Specifically, Janus-flow [203], Mogao [177] and BAGEL [63] adopt a dual-encoder architecture and presents a minimalist architecture that harmonizes AR language models with rectified flow. They decouples the understanding and generation encoders, using SigLIP or the concatenation of SigLIP and SDXL-VAE as the vision encoder for multimodal understanding and SDXL-VAE or FLUX-VAE for image generation. LightFusion [311] reduces the training difficulty of Bagel by initializing independent understanding and generation models. Omni-View [112] further achieves this unification in multi-view image scenarios by adopting a generative paradigm inspired by MVAR [111]. EMMA [105] concatenates the understanding tokens and generation tokens along the channel dimension rather than the token-wise concatenation [63], enabling the model to handle understanding and generation tasks simultaneously. Meanwhile, EMMA replaces the understanding encoder SigLIP2 with a mixture-of-experts architecture to better handle diverse types of input images. Unlike Bagel, which connects generation and understanding across all layers, HBridge [305] is built upon a pre-trained LLM paired with a diffusion-based generative expert, connecting them only via a selective mid-layer bridge. However, the pseudo hybrid encoding design limits the model’s ability to simultaneously leverage both semantic and pixel-level features during generation, as only the pixel encoder is active in the image synthesis process. This decoupling, while beneficial for modularity and training efficiency, prevents the model from fully exploiting semantic cues during image decoding, potentially weakening fine-grained alignment and multimodal compositionality in generative tasks.
@@ 3.4 Any-to-Any Multimodal Models
A recent and notable addition to the any-to-any unified multimodal models is M2-omni [100], which introduces a highly versatile architecture capable of processing and generating a wide variety of modalities, including text, image, video, and audio. M2-omini takes a step forward by incorporating multiple modality-specific tokenizers and decoders, each carefully designed to handle the unique characteristics of different data types. Specifically, it utilizes NaViT [61] to encode videos and images of arbitrary resolution, and combines a pre-trained SD-3 [224] as the image decoder. For audio, M2-omini introduces paraformer-zh [82] to extract audio tokens, and feeds the predicted discrete audio tokens into the pretrained CosyVoice [69] flow matching and vocoder model to generate audio streams. This integration ensures that M2-omini can effectively generate high-quality images, and audio streams from various inputs, making it a truly multi-modal powerhouse. Ming-Omni [2] adheres to the integrated MoE architecture, wherein modality-specific routing is facilitated through dedicated mechanisms tailored for each token, thereby enabling customized routing distributions. To address the multi-scale phenomenon inherent in visual generation [284], Ming-Omni employs multi-scale learnable queries, directed by an alignment strategy, to iteratively generate images progressing from coarse to fine detail. Furthermore, Ming-Omni integrates the audio modality and implements a dual-stage training strategy to mitigate the mutual influence between audio comprehension and generation tasks. The initial stage emphasizes comprehension capabilities, while the subsequent stage concentrates on enhancing generation quality. BLIP3-o [35] also employs learnable queries to bridge multimodal understanding and generation. However, it utilizes two diffusion models: one for learning CLIP embeddings and the other for using CLIP as a condition to generate images. It reveals that flow matching loss is more effective than MSE loss, enabling more diverse image sampling and yielding better image quality. By using Ling-Flash-2.0 [127] as the main backbone, Ming-Falsh-Omni [3] achieves a favorable trade-off between performance and efficiency. Addressing the specific demands of real-time interaction, Qwen3-Omni [337] proposes a Thinker-Talker MoE architecture, which decouples deep reasoning from streaming speech generation to minimize latency without sacrificing intelligence. Furthermore, LongCat-Flash-Omni [281] extends these capabilities into long-context scenarios (up to 128k tokens), utilizing a Shortcut-connected MoE and curriculum learning strategies to effectively model long-range dependencies in video and audio sequences.
@@ 5.3 Evaluation on Interleaved Generation
5.4 Evaluation on Unification
The above evaluation paradigms mainly assess understanding and generation capabilities in isolation, which is insufficient to determine whether unified models can achieve a complementary interaction between understanding and generation tasks. For example, unified models may leverage their understanding ability to enhance generation, or use generative simulation to support more complete understanding.
To fill this critical gap, RealUnify [257] builds the evaluation on unification around two core aspects: (1) Understanding-Enhanced Generation (UEG), which requires reasoning (e.g., commonsense, logic) to guide image generation; and (2) Generation-Enhanced Understanding (GEU), which requires mental simulation or reconstruction (e.g., of transformed or disordered visual inputs) to solve reasoning tasks. RealUnify contains 1,000 carefully human-annotated instances, covering 10 categories and 32 subtasks. This benchmark combines direct end-to-end evaluation with diagnostic step-by-step evaluation, decomposing tasks into separate understanding and generation stages, enabling us to precisely determine whether performance bottlenecks arise from deficiencies in core capabilities or from failures to effectively integrate these capabilities.